OpenStudy (0487308):

WILL GIVE MEDAL! Problem below:

OpenStudy (13ziar):

where

OpenStudy (13ziar):

?

OpenStudy (0487308):

\[\frac{\frac{ x+6 }{ 12 }}{\frac{x-8}{10}}\]

OpenStudy (0487308):

There it is.

OpenStudy (13ziar):

k

OpenStudy (0487308):

@TrojanPoem

OpenStudy (13ziar):

613x-810

OpenStudy (0487308):

How is that even close to the solution?

OpenStudy (13ziar):

not sure but 1x add 612 = 613

OpenStudy (13ziar):

@Nnesha

OpenStudy (0487308):

This is how I did it, but I'm not sure if it's right: \[\frac{ \frac{ x+6 }{ 12 } }{{ \frac{x-8}{10} }}\]

OpenStudy (13ziar):

whats the question then